The Central Division of National, Inc., has operating income of $16,600 on sales revenue of $179,000. Divisional operating assets are $84,600, and management of National has determined that a minimum return of 13% should be expected from all investments.

Required:

a. Using the DuPont model, calculate The Central Division’s margin, turnover, and ROI. (Do not round intermediate calculations. Round your answers to 2 decimal places.)

Central Division
Margin    %
Turnover turns
ROI %

b. Calculate The Central Division’s residual income.

Residual income
Answer #1
Operating income                                                16,600
Sales revenue                                              179,000
Operating assets                                                84,600
Minimum expected return 13%
Margin= Operating income/Sales revenue
Margin= 16600/179000
Margin= 9.27%
Turnover= Sales revenue/Operating assets
Turnover= 179000/84600
Turnover=                                                    2.12 times
ROI= Turnover * Margin
ROI= 2.12*9.27%
ROI= 19.62%
Residual income= (ROI-minimum expected return)*Operating assets
Residual income= (19.62%-13%)*84600
Residual income= 5602